ROVIBRATIONAL SPECTROSCOPY OF CO_{2}-CS_{2}

\underline{C. DUTTON}, R. EIKEY, R. A. BEAUDET , Department of Chemistry, University of Southern California, Los Angeles, CA 90089-0482.

The absorption spectrum of the weakly bound cluster CO_{2}-CS_{2} has been observed by probing the v_{3} asymmetric stretch of CO_{2} near 2350 cm^{−1}. Clusters were formed by supersonic expansion of a mixture of CS_{2}:CO_{2} in a 2:1 ratio, minimizing higher order clusters, in He as the carrier gas. Etalon, reference and signal spectra were recorded simultaneously using a transient digitizer and a 386 PC. Spectral analysis indicates the dimer has a slipped parallel configuration. The structure will be discussed and compared to ab initio results as well as to other CO_{2} complexes.